The leader of the association of jeepney drivers claims that the average daily take home pay of all jeepney drivers in San Mateo, Rizal is P 800. A random sample of 100 jeepney drivers in San Mateo Rizal was interviewed and the average daily take home pay of these drivers is found to be P 825. Use a 0.05 significance level to find out if the average daily take home pay of all jeepney drivers in San Mateo, Rizal is different from P 800. Assume that the population variance is P16, 928.


6. What is the null hypothesis?

7. What is the alternative hypothesis? 

8. What type of test?

9. What is the level of significance?

10. What is the critical value?

11. What is the appropriate test statistic to be used?



1
Expert's answer
2022-03-28T16:01:45-0400

"\text{6. The null hypothesis is}\\
H_{0}: \mu=800\\

\text{7. The alternative hypothesis is }\\
H_{1}: \mu\neq 800\\
\text{8. The test is two tailed}\\

\text{9. The level of significance is 0.05}\\

\text{10. The critical value is}\\
z_{0.025}=1.96\\

\text{11. The appropriate test statistic is}\\
z=\frac{\bar x-\mu}{\frac{\sigma}{\sqrt{n}}}"